Crystallina spent her entire $ 20 weekly allowance on books and candy last week. This week she did the same, but she spent 50 % more on candy and 12.5 % less on books. How much did she spend on candy this week?
Answer by richwmiller(9143)

12.5 less=100-12.5=87.5%=.875
50% more =1.5
1.5c+.875b=20
b+c=20
b=16
c=4
check
1.5*4+.875*16=20
check
6+14=20